OpenStudy (waheguru):

isolate h in pir^2h=v

jimthompson5910 (jim_thompson5910):

\[\Large \pi r^2 h = v\] \[\Large h = \frac{v}{\pi r^2}\]

OpenStudy (waheguru):

thx

jimthompson5910 (jim_thompson5910):

np